OpenStudy (blueraverwolf):

4sinx=-cos^2x+1 Need help please, I don't understand.

OpenStudy (snowflakelove15):

What are you trying to solve?

OpenStudy (blueraverwolf):

Trying to solve for x

OpenStudy (sshayer):

\[4\sin x+\cos ^2x-1=0\] \[4\sin x+(1-\sin ^2x)-1=0\] \[-\sin ^2x+4\sin x=0,\-sin x(\sin x-4)=0 \] sin x=0, \[x=?\] \[\sin x=4 (rejected)~as ~\left| \sin x \right| \leq 1\]